How many solutions exist for the system of equations?
2x-3y=14
4x-6y=21

Answer by ewatrrr(24785)

 
Hi,
How many solutions exist for the system of equations? NONE, Lines are parallel
2x-3y=14 OR y = (2/3)x -14/3
4x-6y=21 OR y = (2/3)x -21/6 |Slopes are Equal